OpenStudy (anonymous):

what is the general solution for the following differential eqation? (x^4)-((c^4)x)=0,where c is a constant.

OpenStudy (anonymous):

IT'S ANSWER IS X=0 OR X=C^(3/4)

OpenStudy (amistre64):

how does that define a diffy q?

OpenStudy (amistre64):

\[\frac{d^4x}{dt^4}-c^{4}x=0\]might be a better construction

OpenStudy (amistre64):

in which case we would find the roots to the characteristic equation: \[r^4-c^4=0\] \[(r^2-c^2)(r^2+c^2)\] \[(r-c)(r+c) (r^2+c^2)\] \[(r-c)(r+c) (r=0\pm ci)\] therefore: \[x=a_1\ e^{-cx}+a_2\ e^{cx}+a_3\ cos(cx)+a_4\ sin(cx)\]